# Fast sampling and model selection for Bayesian mixture models

### v1

@@ A.1 Reformulation of the algorithm
To prove convergence, we first give an alternate formulation of the algorithm, which would be less efficient in practice but for which the proof is simpler. After proving the correctness of this algorithm we then demonstrate that the algorithm of Section III.2—the one we actually use—is equivalent to this alternate formulation.
@@
If $i$ was the only member of component $r$, so that $r$ is now empty, swap the labels of components $r$ and $k$, then delete the empty component $k$ and decrease $k$ by 1. (If $r=k$ then no swapping of labels is necessary; we simply delete the component.)
Assemble a set of $12k+$ 1 candidate states $\mu=(k,z)$, each of which has a weight $w_{\mu}$ associated with it, as follows. {enumerate*}
Of the $12k+$ 1 states, $k$ of them are states in which $i$ is placed in one of the $k$ current components $s$. The weights associated with each of these candidate states are
$w_{\mu}={N-k\over k}P(k)P(x|k,z^{\prime}),$ (29)
where $P(k)$ is the prior on $k$ as previously and $z^{\prime}$ denotes the component assignments in the new state.
@@
$w_{\mu}={k\over k+1}P(k+1)P(x|k+1,z^{\prime}),$ (30)
where $k$ is the number of components before the new component is added.
@@ A.2 Ergodicity
Proof of convergence for any Markov chain Monte Carlo algorithm requires that the algorithm satisfy two conditions: ergodicity and detailed balance NB99 . The requirement of ergodicity states that the algorithm must be able to reach any state of the system from any other in a finite number of Monte Carlo steps. In our algorithm, the states are pairs $k,z$ of number of components plus component assignment for all observations. For these states ergodicity is trivially satisfied by the algorithm above, since our basic Monte Carlo moves can move any observation to a different component or create or delete components, and a finite series of such moves can reach a state with any value of $k$ and component assignment $z$.
@@ A.3 Detailed balance
To prove that our algorithm satisfies detailed balance we must consider several different cases. First, consider moves that do not change the number of components. There are two types of such moves. The first are trivial moves in which we remove the last member $i$ of a component $r$ and delete the component, but then immediately create a new component $s$, making $i$ its sole member. Detailed balance is simple in this case because the reverse of such a move is just another such move. The probability $P(\mu\to\nu)$ of performing the move is equal to the probability $1/k$ of picking component $r$ times the probability $\pi_{\nu}=w_{\nu}/\sum_{\mu}w_{\mu}$ of choosing the particular move. But the factor $1/k$ is the same in both directions, as are the weights $w_{\mu},w_{\nu}$, since the component assignments $z^{\prime}$ in Eq. (30) are the same before and after the move up to a permutation of the labels, and the denominator $\sum_{\mu}w_{\mu}$ is also the same in both directions, since by construction we consider the same set of candidate states. Thus detailed balance is trivially satisfied in this case.
More complicated are the steps where we move an observation from one component to another (without creating or deleting any components). The probability for the forward move in this case is equal to the probability $1/k$ of choosing a particular component $r$, times the probability $1/n_{r}$ of picking a particular observation $i$, times the probability $\pi_{\nu}=w_{\nu}/\sum_{\mu}w_{\mu}$ of choosing the candidate state $\nu$:
@@
where the primed variable $n_{s}^{\prime}$ denotes the value in state $\nu$. The sums in the denominators of (34) and (35) are once again over the same set of states and so have the same value. Using Eq. (29) for the weights, the ratio of probabilities is then
${P(\mu\to\nu)\over P(\nu\to\mu)}={n_{s}^{\prime}\over n_{r}}{w_{\nu}\over w_{\mu}}={n_{s}+1\over n_{r}}{P(x|k,z^{\prime})\over P(x|k,z)},$ (36)
where $z$ denotes the component assignments in state $\mu$, $z^{\prime}$ denotes the assignments in state $\nu$, and we have made use of $n_{s}^{\prime}=n_{s}+1$.
@@
${P(\nu)\over P(\mu)}$ $={n_{r}^{\prime}!n_{s}^{\prime}!\over n_{r}!n_{s}!}{P(k)P(x|k,z^{\prime})/P(x)\over P(k)P(x|k,z)/P(x)}$
$={n_{s}+1\over n_{r}}{P(x|k,z^{\prime})\over P(x|k,z)},$ (38)
@@
Equation (38) is indeed equal to Eq. (36) and hence detailed balance is established. The only exception to this result is for the special case where an observation is “moved” to the same component it is already in, so that no change of state occurs. However, it is easy to demonstrate that, with the choice of weights in Eq. (29), both sides of Eq. (33) are equal to 1 in this case, so again detailed balance is satisfied.
The last class of moves we need to consider are those that do change the number of components, either deleting a component or adding a new one. For the purposes of our proof, and without loss of generality, let us consider the forward move $\mu\to\nu$ to be the one that deletes a component and the reverse move to be the one that adds a component. Thus, the forward move removes the last observation $i$ from a component $r$ and, since the component is now empty, deletes the component and replaces it with the component that was previously labeled $k$, decreasing $k$ by one in the process. Then $i$ is placed in one of the other current components $s$. The probability of this move is equal to the probability $1/k$ of choosing component $r$ times the probability $\pi_{\nu}=w_{\nu}/\sum_{\mu}w_{\mu}$ of choosing to place $i$ in component $s$:
@@
The reverse move removes observation $i$ from component $s$ and makes it the sole member of a new component $r$. The probability of this move is equal to the probability $1/k^{\prime}$ of choosing $s$ from the $k^{\prime}$ possibilities, times the probability $1/n_{s}^{\prime}$ of picking observation $i$, times the probability $\pi_{\mu}$ of choosing the move that labels the new component as component $r$, where once again the primed variables denote values in state $\nu$. Thus,
@@
$={k-1\over k}(n_{s}+1){(N-k+1)/(k-1)\over(k-1)/k}$
×$\hskip 80.00012pt\times{P(k-1)P(x|k-1,z^{\prime})\over P(k)P(x|k,z)}$
$={N-k+1\over k-1}(n_{s}+1){P(k-1)P(x|k-1,z^{\prime})\over P(k)P(x|k,z)}.$ (41)
@@
${P(\nu)\over P(\mu)}$ $={P(k-1)P(x|k-1,z^{\prime})/P(x)\over P(k)P(x|k,z)/P(x)}$
×$\qquad\times{(k-2)!(N-k+1)!/(N-1)!\over(k-1)!(N-k)!/(N-1)!}{(n_{s}+1)!\over n_{s}!}$
$={N-k+1\over k-1}(n_{s}+1){P(k-1)P(x|k-1,z^{\prime})\over P(k)P(x|k,z)},$ (42)
@@
Finally, we observe that the moves that create a new component with label $s=1\ldots k+1$ are all equivalent up to a label permutation and hence, if we don’t care about such permutations, we can lump them all together and represent them with a single move that creates a new component $k+1$ and carries $k+1$ times the weight, Eq. (30), of each individual move, which gives
$w_{\mu}=kP(k+1)P(x|k+1,z^{\prime}),$ (43)
as in Eq. (14). This is algorithm 2 as described in Section III.2 and the one we use in our calculations. Lumping states together like this means that when the algorithm performs a move followed by its reverse move, we may end up not in the state we started with but in an equivalent state with permuted labels. This has no effect on the division of the observations into components or on our estimate of the number of components, but it saves us some time and complexity in the algorithm.
This demonstrates the correctness of our algorithm 2. For algorithm 1 the steps of the proof are closely similar. The main difference is that the factor of $1/n_{r}$ in Eq. (34) and similar equations is replaced by $1/N$, because we are choosing an observation at random rather than a component. At the same time, however, a compensating factor appears in Eq. (10) and the two factors cancel, establishing detailed balance once again.
@@ A.4 Implementation
A number of points about implementation of algorithm 2 are worth mentioning. It is inefficient to implement the algorithm in terms of the component membership variables $z_{i}$. A better approach is to maintain instead a list in a simple array of the observations in each component, along with a record of the number $n_{r}$ of members of the component. This allows one to choose a random member of a component quickly, as required by the algorithm, and a member can be efficiently added to a component by putting them in the first free element of the array. A member can be removed by overwriting them with the last element.
@@
In practice, most of the computational effort of the program is spent calculating the weights of Eqs. (13) and (14), which can be sped up by using a look-up table of values $1/(m+n\alpha)$ for integer $m,n$. Some Monte Carlo steps also require us to renumber entire components. Rather than renumbering each member of a component separately, which would be slow, we instead simply swap pointers to the memory locations containing the membership lists, which can be done in $\mathrm{O}(1)$ time.
Finally, some computational effort can be saved by delaying the removal of a selected observation from its component. It is straightforward to calculate the weights in Eqs. (13) and (14) as if the observation had been removed, which allows us to decide which move will be performed before making any updates. As shown in Table 1, typically a half or more of all moves are null moves that simply place the observation back in the same class it was just removed from, in which case no updates are necessary at all.
@@
With these precautions, the running time for a single Monte Carlo step of the algorithm can be reduced to $\mathrm{O}(k)$, which is optimal since $k$ weights have to be computed for every step.

### v2

@@ A.1 Reformulation of the algorithm
To prove convergence, we first give an alternate formulation of the algorithm, which would be less efficient in practice but for which the proof is simpler. After proving the correctness of this algorithm we then demonstrate that the algorithm of Section 3—the one we actually use—is equivalent to this alternate formulation.
@@
  3. 3.
    If $i$ was the only member of component $r$, delete component $r$, relabel component $k$ to be the new component $r$, and decrease $k$ by 1. (If $r=k$ then no relabeling is necessary.)
  4. 4.
    Assemble a set of $2k+1$ candidate states, each of which has a weight associated with it, as follows.
    1. (a)
      Of the $2k+1$ states, $k$ of them are states in which $i$ is placed in one of the $k$ current components $s$. The weights associated with each of these candidate states are
      $w_{\mu}={P(x|k,z_{-i},z_{i}=s)\over P(x|k,z_{-i})},$ (37)
      where $z_{-i}$ denotes all component assignments except for $z_{i}$.
    2. (b)
@@
      $w_{\mu}={k^{2}\over(k+1)(N-k)}{P(k+1)P(x|k+1,z_{-i},z_{i}=s)\over P(k)P(x|k,z_{-i})},$ (38)
      where $P(k)$ is the prior on $k$ as previously, $k$ is the number of components before the new component is added, and the swapping of the labels is already incorporated into the values of $z_{-i}$.
  5. 5.
@@ A.2 Ergodicity
A Markov chain Monte Carlo algorithm converges to its target distribution if the algorithm satisfies two conditions: ergodicity and detailed balance [42]. The requirement of ergodicity states that the algorithm must be able to reach any state of the system from any other in a finite number of Monte Carlo steps. In our algorithm, the states are pairs $k,z$ of number of components plus component assignment for all observations. For these states ergodicity is trivially satisfied by the algorithm above, since our basic Monte Carlo moves can move any observation to a different component or create or delete components, and a finite series of such moves can reach a state with any value of $k$ and component assignment $z$.
@@ A.3 Detailed balance
To prove that our algorithm satisfies detailed balance we must consider several cases. First, consider moves that do not change the number of components. There are two types of such moves. The first are trivial moves in which we remove the last member $i$ of a component $r$ and delete the component, but then immediately create a new component $s$, making $i$ its sole member. Detailed balance is simple in this case because the reverse of such a move is the same move again, and hence the probabilities $P(\mu\to\nu)$ and $P(\nu\to\mu)$ are equal, which trivially satisfies detailed balance.
More complicated are the steps where we move an observation from one component to another without creating or deleting any components. The probability $P(\mu\to\nu)$ for the forward move in this case is equal to the probability $1/k$ of choosing a particular component $r$, times the probability $1/n_{r}$ of picking a particular observation $i$ from that component, times the probability $p_{\nu}=w_{\nu}/\sum_{\mu}w_{\mu}$ of choosing the target state $\nu$:
@@
where the primed variable $n_{s}^{\prime}$ denotes the value in state $\nu$. The sums in the denominators of (42) and (43) are over the same set of states and so have the same value, and hence, using Eq. (37) for the weights, the ratio of probabilities is
${P(\mu\to\nu)\over P(\nu\to\mu)}={n_{s}^{\prime}\over n_{r}}{w_{\nu}\over w_{\mu}}={n_{s}+1\over n_{r}}{P(x|k,z_{-i},z_{i}=s)\over P(x|k,z_{-i},z_{i}=r)}.$ (44)
where the remaining factors have canceled and we have made use of $n_{s}^{\prime}=n_{s}+1$.
@@
${P(\nu)\over P(\mu)}={n_{r}^{\prime}!n_{s}^{\prime}!\over n_{r}!n_{s}!}{P(k)P(x|k,z_{-i},z_{i}=s)/P(x)\over P(k)P(x|k,z_{-i},z_{i}=r)/P(x)}={n_{s}+1\over n_{r}}{P(x|k,z_{-i},z_{i}=s)\over P(x|k,z_{-i},z_{i}=r)},$ (46)
@@
Equation (46) is indeed equal to Eq. (44) and hence detailed balance is established. The only exception is for the special case where an observation is “moved” to the same component it is already in, so that no change of state occurs. However, it is easy to demonstrate that, with the choice of weights in Eq. (37), both sides of Eq. (41) are equal to 1 in this case, so again detailed balance is satisfied.
The last class of moves we need to consider are those that do change the number of components, either deleting a component or adding a new one. For the purposes of our proof, and without loss of generality, let us consider the forward move $\mu\to\nu$ to be the one that deletes a component and the reverse move to be the one that adds a component. Thus, the forward move removes the last observation $i$ from component $r$ and, since the component is now empty, deletes the component and replaces it with the component that was previously labeled $k$, then decreases the value of $k$ by one. Then $i$ is placed in one of the other current components $s$. The probability of this move is equal to the probability $1/k$ of choosing component $r$ times the probability $p_{\nu}=w_{\nu}/\sum_{\mu}w_{\mu}$ of choosing to place $i$ in component $s$:
@@
The reverse move removes observation $i$ from component $s$ and makes it the sole member of a new component $r$. The probability of this move is equal to the probability $1/k^{\prime}$ of choosing $s$ from the $k^{\prime}$ possibilities, times the probability $1/n_{s}^{\prime}$ of picking observation $i$, times the probability $p_{\mu}$ of choosing the move that labels the new component as component $r$, where once again the primed variables denote values in state $\nu$. Thus,
@@
$\qquad={k-1\over k}(n_{s}+1){P(x|k-1,z_{-i},z_{i}=s)\over P(x|k-1,z_{-i})}{(N-k+1)k\over(k-1)^{2}}{P(k-1)P(x|k-1,z_{-i})\over P(k)P(x|k,z_{-i},z_{i}=r)}$
$\qquad={N-k+1\over k-1}(n_{s}+1){P(k-1)P(x|k-1,z_{-i},z_{i}=s)\over P(k)P(x|k,z_{-i},z_{i}=r)}.$ (49)
@@
${P(\nu)\over P(\mu)}$ $={P(k-1)P(x|k-1,z_{-i},z_{i}=2)/P(x)\over P(k)P(x|k,z_{-i},z_{i}=r)/P(x)}{(k-2)!(N-k+1)!/(N-1)!\over(k-1)!(N-k)!/(N-1)!}{(n_{s}+1)!\over n_{s}!}$
$={N-k+1\over k-1}(n_{s}+1){P(k-1)P(x|k-1,z_{-i},z_{i}=s)\over P(k)P(x|k,z_{-i},z_{i}=r)},$ (50)
@@
Finally, we observe that the moves that create new components with labels $s=1\ldots k+1$ are all equivalent up to a label permutation and hence, if we don’t care about such permutations, we can lump them all together and represent them with a single move that creates a new component $k+1$ and carries $k+1$ times the weight, Eq. (38), of each individual move, which gives
$w_{\mu}={k^{2}\over N-k}{P(k+1)P(x|k+1,z_{-i},z_{i}=k+1)\over P(k)P(x|k,z_{-i})},$ (51)
as in Eq. (15). This is the algorithm described in Section 3 and the one we use in our calculations. Lumping states together like this means that when the algorithm performs a move followed by its reverse move, we may end up not in the state we started with but in an equivalent state with permuted labels. This has no effect on the division of the observations into components or on our estimate of the number of components, but it saves us some time and complexity in the algorithm.
@@ A.4 Implementation
A number of points about implementation of the algorithm are worth mentioning. First, it is inefficient to implement it in terms of the component membership variables $z_{i}$. A better approach is to maintain instead a list in a simple array of the observations in each component, along with a record of the number $n_{r}$ of members of the component. This allows one to choose a random member of a component quickly, as required by the algorithm, and a member can be efficiently added to a component by putting it in the first free element of the array. A member can be removed by overwriting it with the last element.
@@
Some Monte Carlo steps also require us to renumber entire components. Rather than renumbering each member of a component separately, which would be slow, we instead simply swap pointers to the memory locations containing the membership lists, which can be done in $\mathrm{O}(1)$ time.
@@
Some computational effort can also be saved by delaying the removal of a selected observation from its component. It is straightforward to calculate the weights in Eqs. (14) and (15) as if the observation had been removed, which allows us to decide which move to perform before making any updates. With these precautions, the running time for a single Monte Carlo step of the algorithm can be reduced to $\mathrm{O}(k)$, which is optimal since $k$ weights have to be computed for every step.
Appendix B Additional examples
In this appendix we give additional example applications of the LCA algorithm from Section 4.3 to real-world data sets, including one data set significantly larger than any other we consider.
B.1 Problem behaviors among teenagers
This example is an application to a large behavioral data set and illustrates a case in which our method gives different answers from previous approaches. We examine a data set presented by Li et al. [30], which describes survey results on the presence or absence of six problem behaviors in 6504 US teenagers. A previous analysis of a subset of these data by Collins and Lanza [9] found four latent classes. Figure 9 shows the posterior distribution of the number of classes returned by our Monte Carlo method and, as we can see, four classes is a possible fit in this case, but the MAP estimate is five, and six is also more likely than four. We find a longer run is necessary with this data set than for others we consider, in order to achieve consistent results—we ran for 250 000 Monte Carlo sweeps, plus 25 000 for burn-in. Nonetheless the calculation is fast: Li et al. report that their Gibbs sampling calculation took about half an hour; ours takes less than 90 seconds on roughly comparable hardware.
B.2 Online dating
In this example we demonstrate an application of the method to a much larger data set, which comes from the online dating service OKCupid. Upon creating accounts on this service, users are asked a set of questions about themselves, which the service uses to match them with potential partners. Our data set consists of answers from about 60 000 users to 14 of these questions, which ask about gender, sexual orientation, relationship status, income, offspring, body type, diet, drinking and smoking habits, drug use, zodiac sign, religion, and whether they like cats and dogs.
The analysis of this data set takes considerably longer than any of the previous ones, since one sweep now corresponds to 60 000 Monte Carlo steps and hence takes longer to complete. 25 000 sweeps, for example, takes about nine minutes on the author’s laptop. Figure 10 (main panel) shows the posterior distribution of the number of classes and, as we can see, this data set has a much larger number of classes than any of our others, with the distribution strongly peaked around a MAP estimate of $k=31$ classes. Inset in the figure we show the mutual information scores for the 14 questions, calculated using the method of Section 4.3.4. As before, these scores indicate which observations are most informative about component assignments. The most informative observations in this case are (perhaps unsurprisingly) gender and (more surprisingly) whether the respondents like cats and dogs. Apparently the world really is divided between cat people and dog people. High mutual information scores also go to the questions on smoking, drinking, and drugs (which may well be correlated with one another) and to body type and religion. At the bottom of the list are zodiac sign, which is perhaps not surprising, but also status (single, married, etc.) and sexual orientation. The latter seem more surprising, but we should bear in mind that their low mutual information scores merely indicate that they are not correlated with people’s answers to the other questions. They are saying, for example, that being straight or gay doesn’t affect whether you drink or like dogs.
Appendix C Algorithm for the model with
general $\boldsymbol{\eta}$
The algorithm of Section 3, which is used in our example calculations, assumes a Dirichlet-categorical prior over component assignments with concentration parameter $\eta=1$. The methods of this paper can, however, be extended to general values of $\eta$, for which the prior takes the form given in Eq. (8):
$P(z|k,\eta)={1\over N!}(N-k)\,\mathrm{B}(N-k,k\eta)\prod_{r=1}^{k}n_{r}{\Gamma(n_{r}+\eta-1)\over\Gamma(\eta)},$ (52)
where $\mathrm{B}(x,y)$ is Euler’s beta function. The generalization makes use of a Bortz-Kalos-Lebowitz style continuous-time Monte Carlo dynamics [7, 42] in which we keep track of a real-valued time variable $t$ that measures elapsed time in arbitrary units. For general $\eta$, a step of the algorithm is as follows:
  1. 1.
    Define a set of weights $u_{r}$ for components $r=1\ldots k$ by
    $u_{r}=\biggl\{\begin{array}[]{ll}1&\qquad\mbox{if $n_{r}=1$,}\\ (n_{r}-1)/(n_{r}+\eta-2)&\qquad\mbox{if $n_{r}>1$,}\end{array}$ (53)
    then draw a component $r$ at random from the categorical distribution with probabilities $u_{r}/\sum_{s}u_{s}$. Simultaneously, increase the time variable according to
    $t\to t+{k\over\sum_{s}u_{s}}.$ (54)
  2. 2.
    Choose an observation $i$ uniformly at random from component $r$.
  3. 3.
    Remove $i$ from component $r$.
  4. 4.
    If $i$ was the only member of component $r$, delete component $r$, relabel component $k$ to be the new component $r$, and decrease $k$ by 1. (If $r=k$ then no relabeling is necessary.)
  5. 5.
    Assemble a set of $k+1$ candidate states $\mu=(k,z)$, each of which has a weight $w_{\mu}$ associated with it, as follows.
    1. (a)
      Of the $k+1$ states, $k$ of them are states in which $i$ is placed in one of the $k$ current components $s$. Each of these candidate states has an associated weight
      $w_{\mu}={P(x|k,z_{-i},z_{i}=s)\over P(x|k,z_{-i})},$ (55)
      where as previously $z_{-i}$ denotes the set of all component assignments except for $z_{i}$.
    2. (b)
      $w_{\mu}={k(N-k-1)\,\mathrm{B}(N-k-1,(k+1)\eta)\over(N-k)\,\mathrm{B}(N-k,k\eta)}{P(k+1)P(x|k+1,z_{-i},z_{i}=k+1)\over P(k)P(x|k,z_{-i})},$ (56)
      where $k$ is the number of components before the new component is added.
  6. 6.
    Once the set of target states is assembled, choose one of them $\mu$ with probability $p_{\mu}=w_{\mu}/\sum_{\nu}w_{\nu}$ and update the system to that new state.
Sample states are drawn in the normal manner at uniform intervals, but using time as measured by the time variable $t$, rather than time in Monte Carlo steps. For the conventional choice $\eta=1$, this algorithm reduces to the algorithm of the main paper. (We leave the demonstration as an exercise for the interested reader.)
Proof of the correctness of the algorithm follows similar lines to that of Section A.3, but allowing for the effect of the continuous time variable. For moves that do not change the number of components, merely moving an observation from one component to another, the probability for the forward move from component $r$ to component $s$ is equal to the probability of choosing component $r$, times the probability $1/n_{r}$ of picking the particular observation $i$, times the probability $p_{\nu}=w_{\nu}/\sum_{\mu}w_{\mu}$ of choosing the candidate state $\nu$, which gives
$P(\mu\to\nu)={u_{r}\over\sum_{r}u_{r}}\times{1\over n_{r}}\times{w_{\nu}\over\sum_{\mu}w_{\mu}}.$ (57)
For the backward move the probability is
$P(\nu\to\mu)={v_{s}\over\sum_{s}v_{s}}\times{1\over n_{s}^{\prime}}\times{w_{\mu}\over\sum_{\mu}w_{\mu}},$ (58)
where $v_{s}$ are the weights of Eq. (53) for the backward move. Bearing in mind that $n_{r}$ and $n_{s}^{\prime}$ must both be greater than 1 if the number of components does not change, and making use of Eq. (53) for $u_{r}$ and $v_{s}$, the ratio of forward and backward probabilities is then
${P(\mu\to\nu)\over P(\nu\to\mu)}$ $={u_{r}/\sum_{r}u_{r}\over v_{s}/\sum_{s}v_{s}}{n_{s}^{\prime}\over n_{r}}\,{w_{\nu}\over w_{\mu}}$
$={(n_{r}-1)/n_{r}\over n_{s}/(n_{s}+1)}\biggl({n_{s}+\eta-1\over n_{r}+\eta-2}\biggr){P(x|k,z_{-i},z_{i}=s)\over P(x|k,z_{-i},z_{i}=r)}\,{\sum_{s}v_{s}\over\sum_{r}u_{r}}.$ (59)
Now we compare this to the ratio of the desired equilibrium probabilities of the corresponding states. We have
$P(k,z|x,\eta)={P(x|k,z)P(z|k,\eta)P(k)\over P(x)},$ (60)
and hence, using (52), we have
${P(\nu)\over P(\mu)}$ $={n_{r}^{\prime}\Gamma(n_{r}^{\prime}+\eta-1)n_{s}^{\prime}\Gamma(n_{s}^{\prime}+\eta-1)P(x|k,z_{-i},z_{i}=s)P(k)/P(x)\over n_{r}\Gamma(n_{r}+\eta-1)n_{s}\Gamma(n_{s}+\eta-1)P(x|k,z_{-i},z_{i}=r)P(k)/P(x)}$
$={(n_{r}-1)/n_{r}\over n_{s}/(n_{s}+1)}\,\biggl({n_{s}+\eta-1\over n_{r}+\eta-2}\biggr){P(x|k,z_{-i},z_{i}=s)\over P(x|k,z_{-i},z_{i}=r)}.$ (61)
Substituting this expression into Eq. (59) we have
${P(\mu\to\nu)\over P(\nu\to\mu)}={P(\nu)\,(1/k)\sum_{s}v_{s}\over P(\mu)\,(1/k)\sum_{r}u_{r}},$ (62)
which is similar to the standard detailed balance formula, Eq. (41), except for the factors of $(1/k)\sum_{r}u_{r}$ and $(1/k)\sum_{s}v_{s}$. This means that each state $\mu$ will appear not with its normal probability $P(\mu)$ but with modified probability $P(\mu)(1/k)\sum_{r}u_{r}$. However, the probability of sampling from state $\mu$ is multiplied by the amount of time for which the system remains in that state, which is $k/\sum_{r}u_{r}$, following Eq. (54). So the probability of sampling is precisely proportional to $P(\mu)$, as required.
For Monte Carlo steps that change the number of components, we can once more, without loss of generality, consider the forward move $\mu\to\nu$ to be the one that removes the last observation $i$ from a component $r$ and then deletes the empty component, and the reverse move to be the one that creates a new component. The probability of the forward move is equal to the probability of choosing component $r$ times the probability $p_{\nu}=w_{\nu}/\sum_{\mu}w_{\mu}$ of placing observation $i$ in a particular new component $s$:
$P(\mu\to\nu)={u_{r}\over\sum_{r}u_{r}}\times{w_{\nu}\over\sum_{\mu}w_{\mu}}.$ (63)
The backward move removes observation $i$ from component $s$ and makes it the sole member of a new component $r$. The probability for this move is equal to the probability of choosing component $s$, times the probability $1/n_{s}^{\prime}$ of picking observation $i$, times the probability $p_{\mu}$ of choosing the move that creates the new component. As with the proof in Section A, we will initially assume separate weights for creating new components with each possible component label $1\ldots k^{\prime}+1$ and values
$w_{\mu}$ $={k^{\prime}(N-k^{\prime}-1)\,\mathrm{B}(N-k^{\prime}-1,(k^{\prime}+1)\eta)\over(k^{\prime}+1)(N-k^{\prime})\,\mathrm{B}(N-k^{\prime},k^{\prime}\eta)}{P(k^{\prime}+1)P(x|k^{\prime}+1,z_{-i},z_{i}=r)\over P(k^{\prime})P(x|k^{\prime},z_{-i})}$
$={(k-1)(N-k)\,\mathrm{B}(N-k,k\eta)\over k(N-k+1)\,\mathrm{B}(N-k+1,(k-1)\eta)}{P(k)P(x|k,z_{-i},z_{i}=r)\over P(k-1)P(x|k-1,z_{-i})}.$ (64)
Then we have
$P(\nu\to\mu)={v_{s}\over\sum_{s}v_{s}}\times{1\over n_{s}^{\prime}}\times{w_{\mu}\over\sum_{\mu}w_{\mu}},$ (65)
and the ratio of probabilities is
${P(\mu\to\nu)\over P(\nu\to\mu)}$ $={u_{r}/\sum_{r}u_{r}\over v_{s}/\sum_{s}v_{s}}n_{s}^{\prime}{w_{\nu}\over w_{\mu}}={n_{s}+\eta-1\over n_{s}}{\sum_{s}v_{s}\over\sum_{r}v_{r}}(n_{s}+1)$
$\quad\times{k(N-k+1)\,\mathrm{B}(N-k+1,(k-1)\eta)\over(k-1)(N-k)\,\mathrm{B}(N-k,k\eta)}{P(k-1)P(x|k-1,z_{-i},z_{i}=s)\over P(k)P(x|k,z_{-i},z_{i}=r)}.$ (66)
Meanwhile, the ratio of equilibrium probabilities for states $\mu$ and $\nu$ is
${P(\nu)\over P(\mu)}$ $={(N-k+1)\,\mathrm{B}(N-k+1,(k-1)\eta)\over(N-k)\,\mathrm{B}(N-k,k\eta)}{(n_{s}+1)\Gamma(n_{s}+\eta)\over\Gamma(\eta)}{\Gamma(\eta)^{2}\over\Gamma(\eta)n_{s}\Gamma(n_{s}+\eta-1)}$
$\qquad\times{P(x|k-1,z_{-i},z_{i}=s)P(k-1)/P(x)\over P(x|k,z_{-i},z_{i}=r)P(k)/P(x)}$
$=(n_{s}+\eta-1){n_{s}+1\over n_{s}}{(N-k+1)\,\mathrm{B}(N-k+1,(k-1)\eta)\over(N-k)\,\mathrm{B}(N-k,k\eta)}$
$\qquad\times{P(k-1)P(x|k-1,z_{-i},z_{i}=s)\over P(k)P(x|k,z_{-i},z_{i}=r)}.$ (67)
Comparing with Eq. (66), we find that
${P(\nu)\over P(\mu)}={P(\nu)[1/(k-1)]\sum_{s}v_{s}\over P(\mu)(1/k)\sum_{r}u_{r}}.$ (68)
Hence, once again, we have a modified detailed balance condition that implies that each state $\mu$ will appear with probability $P(\mu)(1/k)\sum_{r}u_{r}$. But the probability of sampling each state is multiplied by the length of time for which the system remains in that state, which is $k/\sum_{r}u_{r}$ following Eq. (54), and hence the probability of sampling is exactly proportional to $P(\mu)$, as required.
The final step of the proof, as in Section A, is to combine the $k+1$ moves that create new components numbered $1\ldots k+1$ into a single move that creates a component numbered $k+1$, with $k+1$ times the probability. The weight for this move is given by Eq. (64) times $k^{\prime}+1$, which gives the expression in Eq. (56). As previously, this can result in a permutation of the labels of the components, but it has no effect on the actual partition of the observations, since the component labels are arbitrary. This completes the proof of correctness for the generalized algorithm.
Implementation of the algorithm is straightforward and the slightly higher level of complexity compared with the case for $\eta=1$ does not impact performance significantly. In particular, note that one can calculate in advance tables of $n/(n+\eta-1)$ and $\mathrm{B}(N-k,k\eta)$ for $n,k=1\ldots N$ and then use them to evaluate Eqs. (53) and (56) with little performance penalty.
Appendix D Data sets
In this appendix we describe the data sets used in our examples.
D.1 Epileptic seizures
These data come from a clinical trial of medications for pediatric epilepsy reported by Wang et al. [61]. The data describe the experiences of a single epileptic child participant in the trial over a 140-day period. The participant was observed under baseline conditions for 28 days, then received monthly infusions of intravenous gammaglobulin, as a potential treatment, for the remainder of the observation period, and the data set consists of the number of seizure episodes experienced on each day of the trial, as recorded in a diary by the child’s parent. The complete data set is included in the paper by Wang et al.
D.2 Candy dispenser
This data set comes from an automated candy dispenser in the reception area outside the author’s office, which is stocked with M&Ms brand candy. The dispenser uses a motorized corkscrew mechanism to dispense a handful of M&Ms upon the push of a button. The data set records the number of candies dispensed on 857 pushes of the button and was gathered by Max Jerdee, Conrad Kosowski, Gabriela Fernandes Martins, and Chethan Prakash. The measurements were near-consecutive, but four measurements had to be discarded because of missing data, leaving a total of 853 for analysis. A copy of the data set is available for download on the web at https://umich.edu/˜mejn/mixture.
Table 3: The 32 questions selected from the CBS/New York Times poll for the data set used in this paper, listed in the order in which they were asked on the survey.